A firm has a retention ratio of 48 percent and a sustainable growth rate of 7.10 percent. The capital intensity ratio is 1.70 and the debt-equity ratio is .83. What is the profit margin?

14.53 percent

7.55 percent

13.81 percent

12.83 percent

8.12 percent
Answer #1

Sustainable Growth Rate = Retention Ratio(ROE)

So,

ROE = 0.071/0.48 = 14.79%

Debt/Equity = 0.83

(Debt + Equity)/Equity = 0.83 + 1

Asset/Equity = 1.83 (Debt + Equity = Asset)

Equity = Asset/1.83

ROE = 0.1479 = Net Profit/Equity

(1.83)Net Profit/Asset = 0.1479

Asset = Net Profit(1.83/0.1479) = Net Profit(12.37)

Capital Intensity Ratio = Total Assets/Sales

1.70 = Net Profit(12.37)/Sales

Net Profit/Sales = 1.70/12.37

Net Profit Margin = 13.81%